# InstructPipe: Building Visual Programming Pipelines with Human Instructions

### v1

@@ 7.3. Instructing LLMs Poses Challenges for Both Novices and, Potentially, Experts
As we explained above, one reason that caused mental workload is that non-experts found it challenging to instruct LLM. More interestingly, we found that even we, the inventors of InstructPipe, failed to write optimal instructions. As mentioned in section 5.2, two authors annotated captions for the pipelines offline, and we observed multiple imperfections, especially for the complex ones. For instance, the two captions of 8(c) are “Describe the image and turn it into a cat image” and “Edit an image by updating the image caption”. Neither caption explicitly describes the detailed pipeline flow clearly, and therefore, all the six evaluation trials (section 5) were incomplete (see 11(a) for one example). The average ratio of user interactions is $45.8\%$, more than twice the average value for our multimodal pipelines ($20.8\%$). To further understand the cause of the failure, another author improved the instruction into “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”.
@@
Figure 12. A comparison of InstructPipe generated by two instructions: (a) “Edit an image by updating the image caption”; (b) “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”. See 8(c) for the complete pipeline.
The resulting pipeline is significantly improved but still not perfect (11(b)). The user only needs to turn “Imagen” into another mode so that it also accepts the input “image” node. Revisiting the improved instruction, we instructed InstructPipe with “generate a cat image based on THE updated caption”, which actually missed the input image.
@@
7.4. Online InstructPipe
In this project, we made the following assumption: visual programming and its LLM support are both offline. Note that InstructPipe does have connection to the Internet, and the “offline” here implies that the system does not dynamically update its node library with online resources: every node is pre-defined by the system developers. This statement shows one major limitation of InstructPipe we observed in the open-ended session of our user evaluation: the 27 nodes covered by InstructPipe are insufficient to facilitate all the creative instructions from participants. For example, P9 prompted InstructPipe using “giving me [an] image person walking [a] dog with webcam of myself”. After some clarification, we found that P9 wanted to generate a video in which P9 is walking a dog by taking the webcam as the input of P10’s visual appearance. Such a pipeline is infeasible in our system because it requires a text-driven video generation model. While an extension of InstructPipe with such a text-to-video model as a primitive node can solve this specific issue, we argue that, in practice, users may request many more nodes for their customized uses. Therefore, it is rather important for researchers to investigate a generalizable solution for issues of this category.
In this project, InstructPipe enables users to prototype an ML pipeline with human instruction, and we focus our exploration on a fixed node library. An extension of InstructPipe to a system that covers a wide range of possible online functions is still under-explored. With a next generation of InstructPipe that can find ML models, define nodes and implement the node dynamically with human instruction, we believe the aforementioned issues (e.g., the one raised by P10) can be effectively addressed. In the past several years, the community has already established ML libraries and API services for various models (e.g., by Hugging Face (HuggingFace, 2022)). Such ecosystems provide crucial resources for AI-oriented visual programming systems to dynamically define their primitive nodes, and thus provide more powerful support for users. We encourage future work to investigate how to build a model selector (similar to the Node Selector in Figure 4) that can intelligently select the correct online API to be called in a node. We believe such “online InstructPipe” would provide an unprecedented user experience in which researchers can brainstorm with the system, and the system will automatically return a pipeline with state-of-the-art ML models to accelerate researchers’ creative workflow.
8. Limitations
As discussed in the main content, we notice multiple limitations of InstructPipe. In our current implementation, InstructPipe focuses on generating the graph structure in the pipeline (section 4.1), and therefore the system is not able to generate the property value of the nodes. In some nodes like “tf-model-runner” that exist in Visual Blocks (but not covered by InstructPipe), such property value plays a critical role that directly defines the functionality of the node. Additionally, when InstructPipe detects an undefined generated node, it directly disposes of the line without leveraging such information for other purposes (section 4.4). In the evaluation, we were not able to conduct a large-scale user study with thousands of user study sessions to understand the performance of InstructPipe with both variations from the pipeline factors and the human factors (section 5). Our quantitative results show that InstructPipe fails to significantly reduce participants’ mental workload (fig. 4). Qualitative results further validate this finding and reveal several issues that cause the mental burden when users use InstructPipe (section 7.2).
We also observe two major limitations of the project in addition to those discussed in the main content of the paper. First, our user study participants only engaged InstructPipe in one hour, and therefore, it is unclear whether users will still frequently use and appreciate InstructPipe supports or similar AI assistants in general in a long-term manner. Additionally, all participants are non-experts. Hence, we cannot verify the effectiveness of InstructPipe for users with other levels of expertise. Second, InstructPipe currently cannot detect harmful data or misuse of AI. We believe such a feature is crucial, especially when future work builds “online InstructPipe”(section 7.4), which greatly enhances the generalizability of ML pipeline prototyping capability. Future work must study effective methods to eliminate potential harmful uses when AI assistants become increasingly open-ended.
@@ 9. Conclusion
In this paper, we introduce InstructPipe, an AI assistant that facilitates users to build ML visual programming pipelines with text instructions. We design and implement InstructPipe by decomposing the task into three modules: a node selection module, a code writer, and a code compiler. Results in our technical and system evaluations demonstrate that InstructPipe provides users’ satisfactory “on-boarding” experience of visual programming systems and allows them to rapidly prototype an idea. We further discuss the issues we observed concerning LLMs in visual programming, related to both human factors and technical implementations. We hope that InstructPipe will engage a diverse community to easily develop creative machine learning pipelines.
@@ B.2. Code Writer
The prompt in Code Writer is dynamic, which is dependent on the nodes chosen in Node Selector. Therefore, we cannot provide all the possible prompts in the supplementary materials. Here, we will focus on providing examples of two detailed node configurations utilized in InstructPipe. Figure 7 shows the structure of the prompt utilized in this LLM stage. Figure 13 provides two examples of node configurations (i.e., “Body segmentation” and “PaLI”) that InstructPipe may chose into the highlighted line(s). Each configuration includes keys of “nodeSpecId” (i.e., node types), “description”, “category” and “examples”. For those nodes that support input and output edges, “inputSpecs” and “outputSpecs” specify the sockets and their corresponding valid data types. For example, the output socket name of “Body segmentation” is “segmentationResult”, and its data type is “masks”. Some nodes (e.g., “Body segmentation”) include recommended node(s) (e.g., “Mask visualizer” for “Body segmentation”), and our configuration also contains such information in the dictionary.
Appendix C User Study Pipelines
Figure 14 and Figure 15 visualize two pipelines we required the participants to complete in our user study. Figure 15 is a multimodal pipeline that allows participants to interact with AR effects in real time. Our technical evaluation shows that InstructPipe can generate this pipeline accurately: the averaged ratio of human interactions = 5.2%. Figure 15 is a text-based pipeline that provides participants with a summary of the news searched from Google. The technical evaluation reveals that InstructPipe cannot generate this pipeline accurately: the averaged ratio of human interactions = 27.8%.
Note that even though InstructPipe may be able to complete the pipeline structure in Figure 15 from users’ instruction, we observed that participants still need to fine-tune their keywords to get an ideal pair of sunglasses. Additionally, the default anchor value is “Face top”, so participants need to use the drop-down menu on the “Virtual sticker” node to change the value to “Eyes”. This further motivates us to use the metric of “Time” in addition to the number of user interactions in our study. Our demo video also covers the workflows of these two pipelines.
Appendix D Semi-structured Interview Script

### v2

@@ 7.3. Instructing LLMs Poses Challenges for Both Novices and, Potentially, Experts
As we discussed above, non-experts found it challenging to instruct LLM. More interestingly, we found that even we, the inventors of InstructPipe, failed to write optimal instructions. For instance, the two captions of 13(c) are “Describe the image and turn it into a cat image” and “Edit an image by updating the image caption”. Neither caption explicitly describes the detailed pipeline flow clearly, and therefore, all the six evaluation trials (section 5) were incomplete (see 9(a) for one example). The average ratio of user interactions is $45.8\%$, more than twice the average value for our multimodal pipelines ($20.8\%$). To further understand the cause of the failure, another author improved the instruction into “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”.
@@
Figure 9. A comparison of InstructPipe generated by two instructions: (a) “Edit an image by updating the image caption”; (b) “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”. See 13(c) for the complete pipeline.
The resulting pipeline is significantly improved but still not perfect (9(b)). The user only needs to turn “Imagen” into another mode so that it also accepts the input “image” node. Revisiting the improved instruction, we instructed InstructPipe with “generate a cat image based on THE updated caption”, which actually missed the input image.
@@
8. Limitations and Future Work
8.1. Assisting Humans to Prompt AI Copilot in Visual Programming
InstructPipe introduces a novel user interaction technique for visual programming, along with its set of challenges – prompting AI is not easy (Zamfirescu-Pereira et al., 2023). While latest research has explored prompt writing assistants (Brade et al., 2023; LangChain, 2024), creating such assistants in visual programming poses unique challenges, as discussed in section 7.2, and requires further dedicated investigation. Despite these challenges, the visual programming workspace offers unique opportunity – it provides an interactive and visual medium for delivering AI-generated information. For example, a prompt writing assistant could provide “a pipeline preview” in real time via a lightweight LLM. Visualizing estimated outcomes, such as unexpected pipeline results (as illustrated in 9(a)), highlights the potential of these tools to guide users in refining their instructions effectively.
8.2. Node Parameter Tuning
InstructPipe focuses on generating the graph structure in the pipeline (section 4.1), and InstructPipe is not able to generate node parameters. Latest research in AI agents show great potentials of distributing a systematic task among multiple LLMs and let them solve the problem collaboratively (Jin et al., 2024). We encourage future work to extend such distributed AI agent approaches to generate suitable node parameters to further reduce users’ workload in tuning them.
8.3. A Larger and Dynamic Node Library
InstructPipe is a AI assistant prototype on a small scale library with 27 nodes. Similar to other tool-calling LLM systems (Schick et al., 2024; De La Torre et al., 2024), InstructPipe cannot generate any out-of-scope node, and thus there is a limited scope of pipelines that InstructPipe can generate. Future work should investigate a scale-up problem by creating an assistant that support a large-scale nodes (Qin et al., 2023a). What new technical challenge will emerge? Will such a large-scale library provide practical human value? If yes, what are the concrete new user experience it opens up in visual programming?
Additionally, future work should explore a dynamic solution of the node library, in which an AI assistant can define necessary nodes in visual programming on the fly. HuggingGPT (Shen et al., 2024) is a pioneering project that shares similar vision as this goal, but existing investigations show that the accuracy of such open-ended generation is still unsatisfactory (Patil et al., 2023; Qin et al., 2023b). How can we design an interface to bridge such imperfect AI and human users in visual programming copilot? What will be the interaction paradigm in an interactive system with a dynamic node library?
8.4. Refining System Component Design
InstructPipe provides a system contribution, and we verified the usefulness of InstructPipe via two evaluations that assess InstructPipe as a whole system. One important future direction would be to verify (or even challenge) each technical component of our system, as elaborated below:
Pseudocode. We designed the pseudocode order based on how algorithm papers present their algorithm blocks. Is this design the best approach among all the possible candidates? If not, how can we further improve the design of pseudocode language?
Prompt Design. We leveraged in-context learning capability of LLMs in our prompt design. Prior work shows that few-shot examples cause bias effects in practice (Zhao et al., 2021b), and thus we encourage future work to mitigate this bias by collecting a large dataset and finetuning LLM on the dataset.
Divide-and-conquer at Scale. We adopt the strategy of divide-and-conquer (Smith, 1985) with a two-stage LLM pipeline. Despite its effectiveness with a small node library and simple graphs, its effectiveness is unknown when generating complex graphs. Exploring agent-based approaches (Huang et al., 2023a; Islam et al., 2024) with Retrieval Augmented Generation (RAG) would be a promising future direction to manage complex graph generation in a divide-and-conquer manner (Smith, 1985). We encourage future work to contribute high-quality datasets as well as an interactive LLM system with RAG that provides users with better experiences from AI agents.
8.5. Evaluation Metrics and Long-term Evaluation
In the technical evaluation, we assessed the performance of AI assistants based on the number of user interactions. Existing related metrics, predominantly from the code synthesis literature  (Amini et al., 2019; Hendrycks et al., 2021), largely focuses on two categories: correctness-based metrics  (Chen et al., 2021; Austin et al., 2021) that rely on test case verification, and similarity-based metrics (Tran et al., 2019). Very little research falls outside these two categories (Paul et al., 2024). Our metric incorporates human factors by objectively estimating user effort through graph theory, addressing a gap in the visual programming literature where human-centric considerations are crucial. While our work advances metric development in this domain, further formal research is essential to establish comprehensive standards for visual programming evaluation.
In the user evaluation, we conducted a lab study to understand the user experience of InstructPipe. As future work, we plan to conduct longer-term studies and gather more realistic insights than those we obtained from the lab study. This is critical for us to understand the long-term usefulness of our assistant for beginners, as well as collecting feedback to inform our system design.
8.6. Responsible AI
InstructPipe currently cannot detect harmful data or misuse of AI. We believe such safety features are crucial, especially in the context of the potential for future dynamic node libraries, which would greatly enhance the generalizability of ML pipeline prototyping capability. Future work must study effective methods to eliminate potential harmful uses when AI assistants become increasingly open-ended.
@@ 9. Conclusion
This paper introduces InstructPipe, an AI assistant that empowers users to accelerate their design of ML visual programming pipelines using text instructions. We design and implement InstructPipe by decomposing the task into three modules: a node selection module, a code writer, and a code compiler. Results in our technical and system evaluations suggest that InstructPipe provides users’ satisfactory “on-boarding” experience of visual programming systems and allows them to rapidly prototype an idea, potentially with more than 5X fewer interactions. We further discuss the issues we observed concerning LLMs in visual programming, related to both human factors and technical implementations. We hope that InstructPipe will inspire the community to continue work in accelerated human-AI collaboration for increased expressivity and creativity, for machine learning pipelines, and beyond.
@@ A.3. Processor Nodes
Appendix B System Implementation
B.1. System Prompts Used in LLM Modules
@@ B.1.2. Code Writer
The prompt in Code Writer is dynamic, which is dependent on the nodes chosen in Node Selector. Therefore, we cannot provide all the possible prompts in the supplementary materials. Here, we will focus on providing examples of two detailed node configurations utilized in InstructPipe. Fig. 6 shows the structure of the prompt utilized in this LLM stage. Fig. 10 provides two examples of node configurations (i.e., “Body segmentation” and “PaLI”) that InstructPipe may chose into the highlighted line(s). Each configuration includes keys of “nodeSpecId” (i.e., node types), “description”, “category” and “examples”. For those nodes that support input and output edges, “inputSpecs” and “outputSpecs” specify the sockets and their corresponding valid data types. For example, the output socket name of “Body segmentation” is “segmentationResult”, and its data type is “masks”. Some nodes (e.g., “Body segmentation”) include recommended node(s) (e.g., “Mask visualizer” for “Body segmentation”), and our configuration also contains such information in the dictionary.
((a)) Before layout optimization.
((b)) After layout optimization.
Figure 11. A comparison of the same generated pipeline before and after layout optimization.
B.2. Code Interpreter
Here we provide more low-level implementation details on Code Interpreter. The Code Interpreter parses generated pseudocode into a visual programming pipeline for visualization at the Visual Blocks workspace. Fig. 12 shows the data type definition of graphs, nodes, and edges in the system. The example JSON file to be parsed into the Typescript interface is available at the Visual Blocks website [^5] https://visualblocks.withgoogle.com/. The JSON file is available for data structure exploration by 1) entering an example project and 2) clicking on the “Export” button on the top-right corner.. Our code defines a visual programming pipeline into an array of serialized nodes, $G(N)$. When the user adds a new node to an pipeline, it is adding a new “SerializedNode”, where it contains the edge definition between this new node and other nodes in the current workspace, to the current “SerialedGraph”. This mechanism ensures that nodes can be incrementally added in the order they appear in the pseudocode order while maintaining the integrity of the graph by clearly defining dependencies and data flow between nodes. Algorithm 1 further shows how InstructPipe parses code and incrementally add nodes to formulate a final serialized graph.
{minted}
[ framesep=2mm, baselinestretch=1.2, bgcolor=LightGray, fontsize=, ] typescript /** A serialized graph. */ export declare interface SerializedGraph nodes: SerializedNode[];
/** other properties */
/** A serialized node. */ export declare interface SerializedNode /** The id of the node, e.g., pali_1. */ id: string;
/** The node spec id, e.g., pali. */ nodeSpecId: string;
/** Serialized incoming edges * that connect to this node. */ incomingEdges?: [inputId: string]: SerializedIncomingEdge[] ;
/** other properties */
/** A serialized incoming edge. */ export declare interface SerializedIncomingEdge /** The id of the source node. */ sourceNodeId: string;
/** The id of the output in the source node. */ outputId: string;
Figure 12. The definition of a graph, a node and an edge in the system using the Typescript language. Only core properties of graphic structure definition are presented in this figure.
1
2Input: $C$: the generated texts (i.e., pseudocode) in the string format.
3 Output: $G(N)$: a visual programming pipeline ($SerializedGraph$) that mainly stores an array of $SeirializedNode$ (Fig. 12).
Variables: $T$: a dictionary of parsed tokens that contains $output\_variable\_id$, $node\_id$, $node\_type$, $node\_arguments$; $e$: the incoming edges of a new node, in the format of $SerializedIncomingEdge$; $n$: a new node in the format of $SerializedNode$. $G:SerializedNode[]=[]$ // Initialize $G$ as an empty array
$lines=line\_parser(C)$ // Parse $C$ into lines of code with no pseudocode order changed.
4 for $line$ in $lines$ do
       /* Example: */
      /* $pali\_1\_out=pali\_1:pali(image=input\_image\_1,prompt=input\_text\_1)$/
[…]
5       $T=tokenizer(line)$
       $e:incomingEdges=create\_incoming\_edges(T)$ // create incoming edges for the new node
       $n:SerializedNode=create\_node(T,e)$ // create a new SerializedNode with the incomingEdges and the parsed dictionary
       $G.add\_serializednode(n)$ // add the new SerializedNode to the graph
6
$G=optim\_layout(G)$ // Perform the UI layout optimization, as shown in Fig. 11
7 return $G$
Algorithm 1 Code Interpreter
Appendix C Technical Evaluation
Figure 13. Example pipelines participants built in the workshops.
C.1. Data Post-Processing
  • Isolated graphs. There exist pipelines that include at least one isolated subgraph. The isolated subgraph, as opposed to the main graph, is defined as a graph (or a node) that has no connection to the main graph in the pipeline that provides the main functionality of the pipeline (e.g., the “Image viewer” node on the bottom-left corner of 11(b)). We observed that some participants typically would like to explore the system by working on a separate sub-space. While we acknowledge its usefulness, leaving such “redundant” graphs in the raw data for the evaluation would cause issues when we calculate the number of user interactions (i.e., the metric used in the evaluation that will be defined in the next subsection).
The observation motivated us to post-process the raw data to present more rigorous evaluation results. We first removed incomplete pipelines and the isolated graphs in each pipeline (if there are any). To further enhance the annotation quality, two authors individually annotated the caption of each pipeline separately by referring to the original captions and pipelines authored by the participants. It is important to note that we finished the workshop and the data annotation task before we completed the system implementation. The two authors had no experience using InstructPipe before completing the annotation. We believed this process could effectively enhance the quality of the captions while maintain fairness of the technical evaluation.
As we clarified in section 5.1, the workshop is designed to be an open-ended creation process. This indicates that the dataset inevitably includes out-of-scope nodes like “custom scripts” (in which the participants write code to process the input data and return custom outputs; see 13(b) for an example) and “TFLite model runner” (which call a custom TensorFlow model with a URL input of the model in the TF-Hub).
We removed the pipelines that contain node(s) out of our focus 27 nodes, and selected all the remaining pipelines as our final evaluation set. We argue that this post-processing is critical for reporting a fair accuracy value since InstructPipe can only generate pipelines based on its known node library. The final 48 pipelines (out of 64 pipelines) are comprised of 23 language pipelines, seven visual pipelines, and 18 multi-modal pipelines. Fig. 13 shows three pipelines created by the participants. 13(b) is an example of the pipelines that include out-of-scope nodes, and therefore is not included in the final 48 pipelines. In the technical evaluation, we ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions from two authors for each pipelines) and evaluated the generation results using the metric that will be introduced below.
C.2. Evaluation Metric: The Number of User Interactions
Our definition of the number of user interactions has two important implications. First, a complete pipeline after user interaction does not need to be the same as the corresponding pipeline in the dataset. As long as it fulfills the task described in the caption, we consider the pipeline complete. Second, our definition does not consider interactions of modifying the node parameters, e.g., typing in a text box or selecting a value in a drop-down box. We argue that such interactions are highly node-dependent and are hard to quantify objectively. More importantly, as we explain in section 4.1, the generation of node parameters is out of the scope of this work.
In the technical evaluation with various pipelines, it is unfair to report an averaged absolute value of user interactions because the complexity of the pipelines varies dramatically. For instance, the user may need to make three edits based on a generated result to complete a large pipeline that requires 20 edits from scratch. In another pipeline, the user also needs to do three edits starting from the generated result, but the whole pipeline only takes three edits to finish. Averaging these absolute values does not provide reasonable insights into how accurate the generation is. Therefore, we reported an averaged ratio of user interactions required to complete a pipeline “from our generated pipeline” to that “from scratch” as our target metric in the technical evaluation.
Appendix D User Evaluation
D.1. Semi-structured Interview Script
@@ 6. Which kinds of technologies would be interesting to add?
D.2. User Study Pipelines
Fig. 14 and Fig. 15 visualize two pipelines we required the participants to complete in our user study. Fig. 15 is a multimodal pipeline that allows participants to interact with AR effects in real time. Our technical evaluation shows that InstructPipe can generate this pipeline accurately: the averaged ratio of human interactions = 5.2%. Fig. 15 is a text-based pipeline that provides participants with a summary of the news searched from Google. The technical evaluation reveals that InstructPipe cannot generate this pipeline accurately without further human interaction, and the averaged ratio of additional human interactions is 27.8%. While the generated diagram (with error) is not deterministic, we observed that InstructPipe commonly generates the pipeline in Figure 14 without “URL to HTML” or “PaLM Text Generator” nodes. The error implies that the LLM may misinterpret 1) the data from the “selected text” port of the “String picker” node is the texts on the web instead of the web URL and 2) that “Text processor” has the LLM capability to process the texts instead of simply combining two texts.
Note that even though InstructPipe may be able to complete the pipeline structure in Fig. 15 from users’ instruction, we observed that participants still need to fine-tune their keywords to get an ideal pair of sunglasses. Additionally, the default anchor value is “Face top”, so participants need to use the drop-down menu on the “Virtual sticker” node to change the value to “Eyes”. This further motivates us to use the metric of “Time” in addition to the number of user interactions in our study. Our demo video also covers the workflows of these two pipelines.
D.3. Assistant Provided to the Participants in the User Evaluation
In the user evaluation, our goal is to make the interface condition (either InstructPipe or Visual Blocks) as the only independent variable that changes our dependent variables (section 5.2). Similar to user evaluations of other early-stage HCI research, we had to improvise for lacking system affordances. As an example, we would include help menus and error recovery modals in the future versions of our system but at this early stage we relied on in-person help to nudge and assist our user study participants. We took actions (i.e., assistants) in the user evaluation to ensure the study is under an appropriate amount of control as well as maintain the fairness of our study.
Here we elaborate two examples of assistants we provided in the user study.
In the InstructPipe condition, one participant started their “instructions” by dragging a text box into the visual programming workspace and began typing. When noticing this issue, we kindly asked the participant whether s/he wanted to write instructions or build a pipeline from scratch. S/he then noticed this issue and clicked on the “InstructPipe” button to write prompts. Note that we explicitly taught every participant how to use InstructPipe and asked participants themselves to go through the instruction processes in the training task (Fig. 7).
In the Visual Blocks condition, one participant first dragged a “Virtual sticker” into the workspace when s/he wants to build the multimodal pipeline as required (Fig. 15). After a while, s/he asked us for the meaning of “landmarks” on the first input port of the “Virtual sticker” node (Fig. 15). We then answered this question and provided a hint on the “Face landmark” node (Fig. 15) that could produce the “Face landmarks” required by the “Virtual sticker”. While we had explained all the nodes that the participants need to use in the study in the training task (Fig. 7), we consider such technical questions as reasonable because all of our participants are non-experts. Programming itself is a difficult skill, and it is quite common that people may forget some of the knowledge that they have just learned. Instead of being silent and keeping the participants stuck on a technical issue, we believed offering technical help is an important action we must take to ensure the data quality we collected in the study.
These anecdotes in the user evaluation reveal several limitations of the visual programming system: some designs may not be very intuitive to non-experts. Since the goal of our user evaluation is understanding the benefits of InstructPipe compared to Visual Blocks (without AI assistants), we made our best efforts to take actions to prevent the effects caused by other factors from influencing our data. Meanwhile, we also encourage future work to further explore the system design so that future users can more easily use our assistant in visual programming.
Table 4. The counterbalance sheet of the user evaluation. Each cell is in the format of “Interface / Pipeline”. “Instruct” and “VB” mean the “InstructPipe” and “Visual Blocks” conditions, respectively. “Search” and “Tryon” represent the “text-based pipeline” (Fig. 14) and the “real-time multimodal pipeline” (Fig. 15), respectively.
 ID Step 1 Step 2 Step 3 Step 4
P1 Instruct / Tryon Instruct / Search VB / Tryon VB / Search
P2 VB / Tryon VB / Search Instruct / Tryon Instruct / Search
P3 Instruct / Search Instruct / Tryon VB / Search VB / Tryon
P4 VB / Search VB / Tryon Instruct / Search Instruct / Tryon
P5 VB / Tryon VB / Search Instruct / Tryon Instruct / Search
P6 Instruct / Tryon Instruct / Search VB / Tryon VB / Search
P7 VB / Search VB / Tryon Instruct / Search Instruct / Tryon
P8 Instruct / Search Instruct / Tryon VB / Search VB / Tryon
P9 Instruct / Tryon Instruct / Search VB / Tryon VB / Search
P10 VB / Tryon VB / Search Instruct / Tryon Instruct / Search
P11 Instruct / Search Instruct / Tryon VB / Search VB / Tryon
P12 VB / Search VB / Tryon Instruct / Search Instruct / Tryon
P13 VB / Tryon VB / Search Instruct / Tryon Instruct / Search
P14 Instruct / Tryon Instruct / Search VB / Tryon VB / Search
P15 VB / Search VB / Tryon Instruct / Search Instruct / Tryon
P16 Instruct / Search Instruct / Tryon VB / Search VB / Tryon
D.4. Counter-Balancing and The Replication Number
Table 4 presents how we perform counterbalance in the user evaluation. We counterbalanced both the interface factor (“InstructPipe” and “Visual Blocks”) and the pipeline factors to reduce the learning effects. We then replicated the order four times so that we collected multiple data from different participants in each unique study order. This helps strengthen the power of the data we collected in the study. Note that, in the group of P5 - P8, we flipped the orders within P5 and P6 as well as P7 and P8, but this does not cause a difference in the counterbalance process.